Define marginal utility. Provide an argument why to maximize total utility of good x and good y, the consumer should consume until the ratio of marginal utilities over price is the same across both goods. [6 marks]


1
Expert's answer
2021-05-10T14:35:44-0400

Marginal Utility is the added satisfaction or advantage a customer receives from the purchase of an additional product or service unit.

 The law of diminishing marginal utility theory stares that the increment in consumption of good on continuous basis reduces the marginal utility and raises the total utility at decreasing rate. 

If the marginal utility over price ratios are not equal then the satisfaction of the consumers is not maximized. 

MUxPx>MUyPy\frac{MU_x}{P_x }>\frac{MU_y}{P_y}

The consumption of good X must be increased and consumption of good y must decreased to make these ratios equal.

Case-2:

MUxPx<MUyPy\frac{MU_x}{P_x }< \frac{MU_y}{P_y}

The consumption of good Y must be increased and consumption of good X must decreased to make these ratios equal.